OpenStudy (anonymous):

The perimeter of a rectngle is 54cm. The area of the same rectangle is 176cm^2. What are the dimensions of the rectangle?

OpenStudy (jdoe0001):

\(\bf w=width\qquad l=length \\ \quad \\ perimeter=w+w+l+l\implies 2w+2l = 54 \\ \quad \\ area = l\cdot w = 176\) if you solve for "w" in say \(\bf l\cdot w = 176\) what would it be?